OpenStudy (anonymous):

Write the equation of a line whose slope is -2 and passes through (-1,6). Write the answer in slope-intercept form

OpenStudy (anonymous):

the equation of a line in slope intercept form is y=mx+c y=-2x+c then substitute x=-1,y=6 6=-2+c c=8 therefore y=-2x+8

OpenStudy (anonymous):

Here x1 = -1 and y1 = 6 and m= -2 where m is slope Equation of a line having a slope and passing through a point is giving by y - y1 = m(x - x1) y - 6 = -2 [ x - (-1) ] y - 6 = -2 ( x +1 ) y - 6 = -2x - 2 2x + y = 6 - 2 2x+y=4 2x+y-4=0